OpenStudy (anonymous):

3x + 6y = 18 and x + 2y = 6... Ned someone to correct if im wrong.... I think this has no solutions?

OpenStudy (anonymous):

One solution and that is (6,0)

OpenStudy (across):

This has infinitely many solutions.

OpenStudy (anonymous):

Only one solution

OpenStudy (anonymous):

x-2y=6 multiply by 3 & add. 3x-6y=18 3x-6y=18 ------------ 6x=36 x=36/6 x=6 ans. 6-2y=6 -2y=6-6 -2y=0 y=0 ans. Proof: 3*6-6*0=18 18+0=18 18=18 One solution: (6,0).

OpenStudy (across):

Notice that 3x + 6y = 18 <-> 3(x + 2y = 6)

OpenStudy (across):

In other words, both equations are the same line.

OpenStudy (eyust707):

yes those two equations are the same line